OpenStudy (kaylala):

proving identities: (1-3cos-4cos^2)/(sin^2) = (1-4cos)/(1-cos)

OpenStudy (kaylala):

@UnkleRhaukus

OpenStudy (anonymous):

let U=cos x 1-3cos-4cos^2= 1-3U-4U^2=(1-4U)(1+U) ___> 1-3cos-4cos^2=(1-4cos)(1+cos) sin^2 x+cos^2 x=1 _____> sin^2 x=1-cos^2 x (1-3cos-4cos^2)/(sin^2)=(1-4cos)(1+cos)/(1-cos^2 x) =(1-4cos)(1+cos)/(1-cos x)(1+cos x) =(1-4cos)/(1-cos)

OpenStudy (kaylala):

i dont understand @mustafa2014

OpenStudy (***[isuru]***):

\[\frac{ 1 - 3\cos \theta - 4\cos^{2}\theta }{ \sin^{2}\theta } = \frac{ 1 - 4\cos \theta }{ 1 - \cos \theta }\] before we solve this... let's just remember this \[\cos^{2} \theta + \sin^{2} \theta =1\] which means \[\sin^{2}\theta = 1 - \cos^{2}\theta = ( 1- \cos \theta) ( 1 + \cos \theta) - difference \ of \ 2 \ squares \] now lets consider the left hand side.... \[\frac{ 1 - 3\cos \theta - 4\cos^{2} \theta }{ \sin^{2} \theta }\] factorize the upper part \[\frac{ (1- 4\cos \theta)( 1 + \cos \theta) }{ \sin^{2} \theta }\] now expand sin^2 theta using the expression i mention above \[\frac{ (1 - 4\cos \theta)(1 + \cos \theta) }{ (1-\cos \theta)(1 + \cos \theta) }\] now u have the answer.... ancel out the ( 1+ cos theta) part from both numerator ande denominator and u will have \[\frac{ (1 - 4\cos \theta) }{( 1 - \cos \theta) }\] hope this will help ya!! and srry for taking a long time :)